OpenStudy (anonymous):

what is |3x-12| greater than or equal to 12?

OpenStudy (mathstudent55):

An absolute value inequality of this type, |A| >= b where A is an expression in x, and b is a positive number is solved by changing it into two inequalites separated by the word "or" A >= b or A <= -b In this case, the expression is x is 3x - 12, so to solve |3x - 12| >= 12, solve these two inequalites: 3x - 12 >= 12 or 3x - 12 <= -12